OpenStudy (anonymous):

A student collected the following data for a fixed volume of gas: Temperature (⁰C) Pressure (mm of Hg) 10 726 20 750 40 800 70 880 100 960 150 ??? Fill in the missing data point. Show all calculations leading to an answer help

OpenStudy (cuanchi):

1.- convert all the temperature values from C to K K=C+273 2.- divide each of the pressure by the temperature in K 3.- calculate the average of the P/T calculated in (2) 4.- multiply the 150C converted to K by the average of the P/T calculated in (3)

OpenStudy (greatlife44):

Volume is constant

OpenStudy (greatlife44):

You need to use this \[\frac{ P_{1} }{ T_{1}} = \frac{ P_{2} }{ T_{2}}\]

OpenStudy (greatlife44):

as @cuanchi said you must convert the temperature from celcius to kelvin.

OpenStudy (greatlife44):

\[\frac{ P_{1}*T_{2} }{ T_{1}} = P_{2} \]